OpenStudy (anonymous):

Evaluate the algebraic expression for the given value(s) of the variable(s). x^2-8x+y^2;x=-1,y=-4

OpenStudy (anonymous):

\[-4^2-8(-4)+-1^2\]

OpenStudy (anonymous):

\[16+32+1\]

OpenStudy (anonymous):

49

OpenStudy (aum):

\(x^2-8x+y^2 \\ x=-1,~~y=-4 \\ (-1)^2 -8(-1) + (-4)^2 = \\ 1 + 8 + 16 = 25 \)

OpenStudy (anonymous):

Thank you aum!!!

OpenStudy (aum):

You are welcome.